OpenStudy (anonymous):

Evaluate the expression. Assume that all angles are in Quadrant I. sec(cos-1 3/5) a. 3/5 b. 4/5 c. 5/4 d. 5/3

OpenStudy (badhi):

 find sec(a)

OpenStudy (anonymous):

5/3

OpenStudy (anonymous):

but its like a double secant right? so does it switch back to 3/5? @BAdhi

OpenStudy (anonymous):

anybody please help!

OpenStudy (shamim):

ya i wanna try it

OpenStudy (shamim):

let \[\cos^{-1} (3/5)= \theta\]

OpenStudy (shamim):

\[or, \cos \theta =\frac{ 3 }{ 5 }\]

OpenStudy (shamim):

now we can draw a right triangle

OpenStudy (shamim):